Q.1a box experienced a force of 200 N east for 8.7 s. calculate the impulse caused by the force
Q.2a bullet of mass 0.004 kg moving north with an initial velocity of 700 ms-1is subjected to a force of 7N north for 3.5s calculate the final velocity of the bullet
1
Expert's answer
2019-02-07T09:28:07-0500

1) By the definition of the impulse we get:

"J = F \Delta t = 200 N \cdot 8.7 s = 1740 N \cdot s."

2) We can find the final velocity of the bullet from the definition of the impulse:

"J = F \Delta t = \Delta p = m \Delta v = m(v_f - v_i),""F \Delta t = mv_f - mv_i,""v_f = \frac{F \Delta t + mv_i}{m} = \frac{7 N \cdot 3.5 s + 0.004 kg \cdot 700 ms^{-1}}{0.004 kg} = 6825 ms^{-1}."

Answer:

1)

"J = 1740 N \cdot s."

2)

"v_f = 6825 ms^{-1}."